OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0. Enter exponents using the caret (^). 9-5x^2=(x-2)^2-4

OpenStudy (anonymous):

We already answered this... Remember? \[-6 x^2+4 x+9 = 0\]

OpenStudy (anonymous):

i dont know how to do it can you help mee please

OpenStudy (anonymous):

We wrote it on the History one!

OpenStudy (anonymous):

9−5x2=(x−2)2−4 9−5x2=x2−4x −6x2+4x+9=0

OpenStudy (anonymous):

Yes! You move everything to the left side...